### v1

# ReCo-Diff: Explore Retinex-Based Condition Strategy in Diffusion Model for Low-Light Image Enhancement

Abstract
Low-light image enhancement (LLIE) has achieved promising performance by employing conditional diffusion models. Despite the success of some conditional methods, previous methods may neglect the importance of an explicit formulation of task-specific condition strategy, resulting in suboptimal visual outcomes. In this study, we propose ReCo-Diff, a novel approach that incorporates Retinex-based prior as an additional pre-processing condition to regulate the generating capabilities of the diffusion model. ReCo-Diff first leverages a pre-trained decomposition network to produce initial reflectance and illumination maps of the low-light image. Then, an adjustment network is introduced to suppress the noise in the reflectance map and brighten the illumination map, thus forming the learned Retinex-based condition. The condition is integrated into a refinement network, implementing Retinex-based conditional modules that offer sufficient guidance at both feature- and image-levels. By treating Retinex theory as a condition, ReCo-Diff presents a unique perspective for establishing an LLIE-specific diffusion model. Extensive experiments validate the rationality and superiority of our ReCo-Diff approach. The code will be made publicly available.
1 Introduction
Figure 1: Visual comparisons among recent DiffLL [14], PyDiff [57], and our ReCo-Diff on an low-light image from the LOL-v2 dataset. Previous diffusion-based methods exhibit detail loss and color distortion. Our method properly maintains color constancy and generates realistic textures thanks to the introduction of the superior Retinex-based condition strategy.
Low-light photography is quite prevalent in the real world due to inherent environmental or technology restrictions, such as inadequate lighting and limited exposure times. Images captured in weakly illuminated conditions not only have low visibility for human perception but also are unsuitable for subsequent multimedia computing and downstream vision tasks, such as object detection [3, 19] and semantic segmentation [6]. Therefore, a large number of methods for low-light image enhancement (LLIE) are proposed to expose hidden features in low-light images and avoid performance degradation in subsequent vision tasks.
Thanks to the development of diffusion models (DMs) [13, 33], numerous diffusion-based studies have been conducted for image restoration tasks [31, 28], with the goal of facilitating texture recovery. Previous diffusion-based image restoration methods choose to concatenate the low/high-quality images [31, 30] and extract features and priors [8, 10] as conditions. Furthermore, conditional DMs have already been introduced in LLIE tasks and thus several successful attempts have emerged [48, 14, 37, 57, 49]. It is worth noting that sampling efficiency is a common difficulty for the diffusion model, where limited sampling steps result in subpar generation fidelity. Therefore, Jiang et al[14] adopt 2D discrete wavelet transformations and utilize the low-resolution coefficients as conditions for faster inference speed. However, without the explicit modeling of color information, the results are unpleasing as shown in Fig. 1(a). To better adapt to LLIE task, the color map [49, 34, 57] and Retinex model [48] are introduced into the diffusion process. [34] and [49] try to maintain color consistency, while they only apply an invariant color map and gain limited ability of enhancement. Zhou et al. [57] propose a pyramid diffusion model (PyDiff) to perform the reverse process by progressively increasing resolution and adjusting the color after denoising through a global corrector, which improves the visual quality to a certain extent as shown in  Fig. 1(b). However, the visual results show limitations in preserving color and details, which may be solved by the solid Retinex theory. Thus, to facilitate the physical interpretability, Diff-Retinex [48] explores the possibility of establishing the Retinex-based diffusion model. The denoising network inputs reflectance and illumination maps as conditional images, and a consistent network is proposed to preserve the content, similar to PyDiff [57]. However, directly applying the decomposed maps as conditions is expected to be imperfect since Retinex theory is essentially an ideal model and the corruptions in reflectance and illumination need to be considered [1]. Although numerous studies have recognized the significance of auxiliary guidance in the diffusion process, they fail to explore a suitable strategy and still produce unfavorable visual results. Thus, the condition strategy has the potential for further improvement in establishing an LLIE-specific diffusion model with significant capability of enhancement.
To address the aforementioned issues, we propose a novel Retinex-conditioned diffusion model, ReCo-Diff, specially designed for the LLIE task. From a new perspective, we combine physical models and the conditioning technique to produce better results. Specifically, our ReCo-Diff learns the Retinex-based prior as extra pre-processing conditions in the diffusion model and refines the enhanced output by integrating the Retinex-based conditions as guidance, instead of processing the reflectance and illumination maps directly [48]. In each step, our ReCo-Diff conducts Retinex-based condition learning and conditional image refinement successively. In the stage of condition learning, we first utilize a pre-trained network (DNet) to produce initial decomposed maps. The adjustment network (ANet) improves the initial maps by suppressing noise in reflectance and adjusting exposure in illumination, and the outputs serve as reliable Retinex-based conditions for better refinement. Then, different from simply using color maps as conditional inputs [49, 57, 34], our ReCo-Diff incorporates the Retinex-based conditions into the denoising U-Net (UNet) for better conditional guidance. Furthermore, unlike [57, 48] directly using stacked convolutional layers to improve the enhanced image, we propose a refinement network (RNet) to better preserve color and contents based on the Retinex theory, which separates the illumination while maintaining the color and details in the reflectance. We reformulate the Retinex model into a residual manner to alleviate the corruptions and thus design the feature- and image-level Retinex-conditioned modules (F/IRCM) in the RNet. The physically inspired conditions and the modified Retinex model for refinement endow our ReCo-Diff with the superior capability of enhancement. Finally, as shown in Fig. 1(c), our ReCo-Diff provides the most pleasing result by using the novel task-specific condition strategy.
The main contributions of our work are as follows:
  • We propose a novel diffusion-based method for LLIE with a Retinex-based condition strategy, exploring the role of physical models in controlling the generation capabilities of diffusion models.
  • We formulate our ReCo-Diff from two aspects: the condition learning stage that improves initial decomposed maps and produces reliable Retinex-based conditions; and the conditional refinement stage that exploits the Retinex-based conditions to guide the diffusion process for better preserving the color and contents.
  • Extensive experiments on representative benchmarks demonstrate that the Retinex-based condition strategy and the well-designed learning and refinement process lead to the superior performance of our ReCo-Diff.
Figure 2: Overview of our Retinex-based conditional diffusion model (ReCo-Diff). Each step contains two stages: (a) In the learning stage, the pre-trained DNet provides the initial decomposed maps and the ANet adjusts and outputs reliable Retinex-based conditions. (b) In the refinement stage, the learned conditions are incorporated into UNet as extra inputs and are incorporated into RNet through F/IRCMs to preserve the color and content. FRCM and IRCM respectively receive the features and adjusted maps from ANet, detailed in Fig. 3.
2 Related Work
Low-light Image Enhancement. Numerous works introduce the Retinex theory into deep neural networks [39, 52, 53, 40, 1, 48, 5, 7, 21, 22]. The Retinex-Net [39] is the most inspiring method combining physical model and DNNs. Then, Zhang et al. proposes KinD [52] and KinD++ [53] to provide more effective solutions. Rather than complex multi-stage training pipeline, Fu et al[5] and Cai et al[1] explore the possibility of end-to-end frameworks and achieve significant performance improvement.
Without Retinex theory, recent works concentrate on directly end-to-end manners [42, 26, 55, 36, 50, 43, 29, 44, 16]. LLNet [27] inspires the emergence of end-to-end methods. Primarily, supervised methods show promising capability of enhancement. For reducing color deviation,  [45, 20, 54, 41] adopt 3DLUT and histogram to preserve color consistency. In [43, 44], the SNR-aware prior and the structure-aware features are taken as guidance to produce realistic results. Recently, the Ultra-High-Definition (UHD) images become popular. LLformer [35] and UHDFour [24] are proposed to enhance UHD images and release UHD datasets to promote the following research. Furthermore, semi-supervised [46, 29], unsupervised [15, 7] and zero-shot learning [9, 56, 23] are valuable when training images are limited.
Diffusion-Based Image Restoration and Low-Light Image Enhancement. Diffusion models [13, 33] show promising capability in image generation tasks [32, 2]. For solving image processing tasks, diffusion-based methods employ conditional mechanism to incorporate the distorted images as guidance [25], such as colorization [31], super-resolution [32], restoration [28, 18, 4], and LLIE [34, 48, 49, 37, 57, 14]. To realize efficient diffusion-based LLIE, DiffLL [14] uses wavelet transformation to decrease the input size and a high-frequency restoration module to maintain the details. PyDiff [57] directly down-samples the image in early steps to speed up the sampling process. Considering the characteristics of LLIE task, LLDiffusion [34] and CLE-Diffusion [49] use the color map as extra conditional input to preserve the color information. Furthermore, PyDiff [57] proposes a global corrector to alleviate color degradation. By introducing Retinex theory, Diff-Retinex [48] decomposes the image and utilizes the reflectance and illumination maps as conditional images to form the guidance.
Although existing diffusion-based LLIE methods obtain good performance, they are limited by insufficient conditional guidance and produce unsatisfactory color and details. In this paper, we investigate a new condition strategy and propose ReCo-Diff to control the generative diffusion model by forming a Retinex-based conditional process, exploring a novel perspective for diffusion-based LLIE and achieving promising performance.
3 Method
The overview is shown in Fig. 2. We present the Retinex-based condition strategy to explore an effective diffusion-based method for LLIE. We introduce the conditional DDPM in Sec. 3.1 and present our ReCo-Diff in Sec. 3.2.
3.1 Conditional Denoising Diffusion Model
To deal with image restoration tasks, condition strategies are proposed to develop the conditional DDPM [25]. The conditional DDPM also generates a target image $\mathbf{x}_{0}$ from a pure noise image $\mathbf{x}_{T}$ and refines the image through successive iterations. Unlike DDPM, the low-quality images are utilized as conditional inputs in conditional DDPM [32]. Thus, the conditional inference is defined as a reverse Markovian process:
$\begin{split}p_{\theta}(\mathbf{x}_{0:T}|\mathbf{y})&=p(\mathbf{x}_{T})\prod\nolimits_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y}),\\ p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y})&=\mathcal{N}(\mathbf{x}_{t-1}\mid\mu_{\theta}({\mathbf{x}}_{t},\mathbf{y},\gamma_{t}),\sigma_{t}^{2}\mathbf{I}),\end{split}$ (1)
where $\gamma_{t}=\prod_{i=1}^{t}\alpha_{i}$ and $\alpha_{1:T}$ are scale parameters. Due to the absence of high-quality image $\mathbf{x}_{0}$ in inference, ${\bm{\epsilon}}_{\theta}$ estimates noise and approximates $\mathbf{x}_{0}$ as:
$\hat{\mathbf{x}}_{0}=\frac{1}{\sqrt{\gamma_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-\gamma_{t}}\,\bm{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{y},\gamma_{t})\right).$ (2)
Therefore, the mean of $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y})$ can be parameterized by applying the $\hat{\mathbf{x}}_{0}$ into the posterior distribution [13]:
$\!\mu_{\theta}(\mathbf{y},{\mathbf{x}}_{t},\gamma_{t})=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\gamma_{t}}}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{y},\gamma_{t})\right).$ (3)
The training objective is to approximate the precise mean $\tilde{\mu}_{\theta}$, which only optimizes the denoising network $\bm{\epsilon}_{\theta}(\cdot)$. The overall optimization objective can be formulated as:
$\mathbb{E}_{\mathbf{x}_{0},\mathbf{y},t,\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\lVert\bm{\epsilon}_{t}-\bm{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{y},\gamma_{t})\rVert^{2}_{2}].\vspace{-0.05cm}$ (4)
3.2 Retinex-Based Conditional Diffusion Model
Retinex theory illustrates the assumption that the reflectance and illumination components can describe the original image, and the reflectance map is consistent under various lighting conditions. Low-/normal-light images $I,\hat{I}\in\mathbb{R}^{W\times H\times 3}$ share the constant $\hat{R}\in\mathbb{R}^{W\times H\times 3}$ and consist of diverse $L,\hat{L}\in\mathbb{R}^{W\times H\times 1}$ as:
$I=\hat{R}\odot L,\quad\hat{I}=\hat{R}\odot\hat{L},\vspace{-0.2cm}$ (5)
where $\odot$ denotes the element-wise multiplication. According to multi-scale Retinex [17], $\hat{R}$ is the Retinex output by removing the lighting effects in $I$ as follows:
$\begin{split}\hat{R}=\log(&I)-\log(L)=\log(I)-\log(\mathcal{G}(I)),\\ \hat{I}&=\mathcal{T}(\exp(\hat{R})),\quad\hat{I}\in(0,1),\end{split}$ (6)
where $\mathcal{G}(\cdot)$ and $\mathcal{T}(\cdot)$ denote the convolution with the Gaussian surround function and the linear transformation function. However, although the ideal model maintains color constancy by estimating the reflectance map $\hat{R}$, details are easily broken through the removal of illumination. Thus, $\hat{R}$ is more suitable to guide the color recovery. We introduce the low-light input to preserve the original information and reformulate Eq. 6 to model the process as:
$\hat{I}=\mathcal{F}(I,\hat{R}),\vspace{-0.18cm}$ (7)
where $\mathcal{F}(\cdot)$ denotes the deep network and $\hat{R}$ acts as an auxiliary guidance. Inspired by the improved Retinex model, we consider the enhancement process from two stages: Retinex-based guidance adjustment and conditional image enhancement. Recalling the conditional DDPM model discussed in Sec. 3.1, the enhancement task can be performed due to the superiority of the generative model conditioned by low-light images [34, 14]. However, the noise and artifacts in low-light images inevitably mislead the diffusion model and cause detail loss and unsatisfactory color.
To mitigate this issue, we utilize the guided enhancement manner in Eq. 7 and integrate the Retinex-based priors into the diffusion process. Notably, color constancy can actually be affected by lighting variations and the corrupted $R$ may lead to degradation of the enhanced image. Thus, we introduce the adjusted term $\hat{\mathbf{c}}=\mathcal{F_{A}}(R,L)$, including more sufficient Retinex-based guidance and acting as priors in the conditional generation process. Subsequently, we formulate the enhancement model as:
$\hat{I}=\mathcal{F_{R}}(\bm{\epsilon}_{\theta}(I,\mathcal{F_{A}}(R,L)),\mathcal{F_{A}}(R,L)),\vspace{-0.18cm}$ (8)
where $\bm{\epsilon}_{\theta}(\cdot)$, $\mathcal{F_{R}}(\cdot)$, and $\mathcal{F_{A}}(\cdot)$ denote UNet, RNet, and ANet in Fig. 2, respectively. Combining the promising texture generation capability of the diffusion model and the conscious representation with vivid color and detail of Retinex theory, we develop a Retinex-based condition strategy for the conditional DDPM and formulate our ReCo-Diff. Eq. 8 represents the denoising and refinement process in one iteration, which can be described as two stages, Retinex-based condition learning and Retinex-conditioned refinement. In the first stage, we obtain the decomposition results from a pre-trained DNet and input them into the following ANet to produce Retienx-based conditions, which will be described in Sec. 3.2.1. Then, the conditions will be introduced into the UNet and RNet to achieve conditional image denoising and refinement as elaborated in Sec. 3.2.2.
3.2.1 Retinex-based Condition Learning
As described in Sec. 3.2, the basis of conditional image refinement is to learn high-quality conditions. Inspired by Retinex-based methods [39, 52], we propose the DNet to serve as a pre-processing module and produce initial Retinex-based conditions. Compared to calculating an invariant color map as conditional input [34, 49], the decomposed maps contain not only color information but also details and illumination, achieving a more sufficient form of conditional inputs. To further exploit the progressively refined images in the diffusion process, DNet inputs $\mathbf{x}_{t}$ as well and provides competent updated predictions $R_{t-1}$ and $L_{t-1}$ for better generation. However, the images in the early steps contain incorrect contents and mislead the decomposition. Therefore, an adjustment network (ANet) is introduced to deal with corruption and ameliorate the low-quality conditions in the early steps, achieving approving conditional enhancement during the whole generative process. Thus, we realize the Retinex-based condition learning as:
$\hat{\mathbf{c}}_{t-1}=\mathcal{F_{A}}(\mathbf{c}_{t-1})=\mathcal{F_{A}}(\mathcal{F_{D}}(\mathbf{y},\mathbf{x}_{t})).$ (9)
Pre-trained DNet. DNet adopts a lightweight UNet-like network to learn the decomposition mapping. It inputs the images $I$ and outputs the reflectance $R$ and illumination $L$. Inspired by the training strategy in [39], we utilize the constant reflectance loss, smooth illumination loss, and reconstruction loss to pre-train DNet. More details can be seen in the supplementary materials.
Condition Adjustment. Directly using the initial decomposed maps with noise and inaccurate brightness may affect the diffusion process. Thus, we utilize the ANet to adjust the decomposition priors. The ANet employs a similar architecture to the DNet with tiny modifications on input and output layers. It inputs the initial maps $\mathbf{c}_{t}=[R_{t},L_{t}]$ and learns to produce the adjusted $\hat{\mathbf{c}}_{t}=[{R_{t}}\!^{\prime},{L_{t}}\!^{\prime}]$ from two aspects. We first aim to suppress the noise and calibrate the color in $R_{t}$ by the joint adjustment loss:
$\mathcal{L}_{ja}=\parallel{R_{t}}\!^{\prime}-\hat{R}\parallel_{1}+SSIM({R_{t}}\!^{\prime},\hat{R}).$ (10)
Then, we propose the joint exposure loss to adjust the sub-optimal illumination as follows:
$\mathcal{L}_{je}=\parallel{L_{t}}\!^{\prime}-\hat{L}\parallel_{1}+Hist({L_{t}}\!^{\prime},\hat{L}),$ (11)
where $Hist(\cdot)$ denotes the L1 loss between the histogram of ${L_{t}}\!^{\prime}$ and ground-truth $\hat{L}$. Thus, the ANet is trained as:
$\mathcal{L}_{ANet}=\lambda_{ja}\mathcal{L}_{ja}+\lambda_{je}\mathcal{L}_{je},$ (12)
where $\mathcal{L}_{ANet}$ is the only loss to propagate gradient for smooth optimization by detaching ${R_{t}}\!^{\prime}$ and ${L_{t}}\!^{\prime}$. Subsequently, ANet learns the progressive adjustment mapping of reflectance and illumination by incorporating time embedding and provides $\hat{\mathbf{c}}_{t}$. Furthermore, the multi-scale features $F_{t}$ are crucial to fully exploit the learned mapping. Thus, the features are included in conditions as $\hat{\mathbf{c}}_{t}=[{R_{t}}\!^{\prime},{L_{t}}\!^{\prime},F_{t}]$, providing better lightness and color guidance.
3.2.2 Retinex-conditioned Refinement
After obtaining the Retinex-based conditions $\hat{\mathbf{c}}_{t}$, the next stage is to incorporate the conditions into the iterative diffusion process. Previous methods propose to introduce the invariant color maps [49, 34, 57] and Retinex decomposition [48]. However, existing condition strategies are insufficient to conduct the favorable iterative enhancement. In our ReCo-Diff, we already learn the Retinex-based conditions as described in Sec. 3.2.1. Then, following the usage of color map [49, 34], we first control the generation capability of UNet by leveraging $\hat{\mathbf{c}}_{t}$ as extra inputs, while the carefully prepared conditions are being underutilized just for denoising. Thus, as illustrated in Eq. 8, we apply a post-processing refinement network (RNet) to fully exploit the conditions and guarantee the consistent color and contents of ${\hat{\mathbf{x}}_{0}}^{\prime}$. Different from [48], RNet adopts a lightweight architecture and implements Retinex-conditioned refinement at both feature- and image-levels by FRCM and IRCM. Overall, the refinement stage can be elaborated as:
${\hat{\mathbf{x}}_{0}}^{\prime}=\mathcal{F_{R}}(\bm{\epsilon}_{\theta}(\mathbf{x}_{t-1},\hat{\mathbf{c}}_{t}),\hat{\mathbf{c}}_{t}),$ (13)
where ${\hat{\mathbf{x}}_{0}}^{\prime}$ denotes the output refined by RNet at time step $t$.
Conditional Denoising. UNet introduces reflectance and illumination maps through embedding layers and integrates them in feature space instead of concatenation for effective information interaction. Specifically, we first transform the original ${R_{t}}\!^{\prime}\in\mathbb{R}^{W\times H\times 3}$ and ${L_{t}}\!^{\prime}\in\mathbb{R}^{W\times H\times 1}$ into embedding features $R_{em}$ and $L_{em}$. To maintain the content and color information, the reflectance embedding $R_{em}\in\mathbb{R}^{W\times H\times C}$ keeps the original resolution. Moreover, the illumination embedding is downsampled as $L_{em}\in\mathbb{R}^{W/2\times H/2\times C}$ for computational efficiency thanks to the smooth distribution of lightness. Then, the $R_{em}$ and $L_{em}$ are input to the conditional layers with original and half sizes respectively, guiding the denoising process. Thus, we can approximate $\hat{\mathbf{x}}_{0}$ by updating Eq. 2 as:
$\hat{\mathbf{x}}_{0}=\frac{1}{\sqrt{\gamma_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-\gamma_{t}}\,\bm{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{y},\hat{\mathbf{c}}_{t},\gamma_{t})\right).\vspace{-0.15cm}$ (14)
Figure 3: The architecture of the feature/image-level Retinex-conditioned modules (F/IRCM). FRCM inputs multi-scale features $F_{t}$ from ANet and obtains the optimized image feature $F_{\hat{\mathbf{x}}_{0}}^{\>{}^{\prime}}$. Then, IRCM inputs the ${R_{t}}\!^{\prime}$, ${L_{t}}\!^{\prime}$ and the $F_{\hat{\mathbf{x}}_{0}}^{\>{}^{\prime}}$ to refine the approximated $\hat{\mathbf{x}}_{0}$ a produce the final output ${\hat{\mathbf{x}}_{0}}^{\prime}$.
Conditional Refinement. As shown in Fig. 2, RNet consists of two Retinex-conditioned modules to realize the color recovery and detail enhancement of the approximated $\hat{\mathbf{x}}_{0}$. Specifically, we first implement the Retinex-conditioned refinement in feature space, since $F_{t}$ incorporates rich information of adjustment mapping as illustrated in Sec. 3.2.1. Thus, as depicted in Fig. 3, we unify the sizes of multi-scale features in $F_{t}$ and model them as affine transformation parameters $F_{\gamma}$ and $F_{\beta}$. Then, the feature fusion is carried out by scaling and shifting operations $\mathcal{F}_{T}(\cdot|\cdot)$ as:
$F_{\hat{\mathbf{x}}_{0}}^{\>{}^{\prime}}=\mathcal{F}_{T}(F_{\hat{\mathbf{x}}_{0}}|F_{t})=F_{\gamma}\odot W(F_{\hat{\mathbf{x}}_{0}})+F_{\beta},\vspace{-0.15cm}$ (15)
where $W$ denotes the convolution layer and $\odot$ is the dot product. However, although the FRCM and conditional layers in UNet integrate the Retinex-based priors into the diffusion process, applying the calculation of the superior Retinex model can further refine the enhanced result. Therefore, as shown in Fig. 3, we conduct refinement at image-level by introducing ${R_{t}}\!^{\prime}$ and ${L_{t}}\!^{\prime}$ into IRCM. To simultaneously alleviate the errors resulting from the ideal model and establish a stable training process, we reformulate Eq. 6 into a residual refinement manner as:
$\begin{split}\Delta\hat{\mathbf{x}}_{0}&=\mathcal{F}_{T}({R_{t}}\!^{\prime}|\mathcal{F}_{T}({L_{t}}\!^{\prime}|\hat{\mathbf{x}}_{0})),\\ {\hat{\mathbf{x}}_{0}}^{\prime}&=\mathcal{F}_{T}(\hat{\mathbf{x}}_{0}|\Delta\hat{\mathbf{x}}_{0})+W(F_{\hat{\mathbf{x}}_{0}}^{\>{}^{\prime}}),\end{split}$ (16)
where ${\hat{\mathbf{x}}_{0}}^{\prime}$ is the final refined output of RNet. In RNet, FRCM and IRCM fully exploit the Retinex-based conditions and produce results with consistent color and details. Accordingly, in addition to Eq. 4, we also utilize the constraint of ${\hat{\mathbf{x}}_{0}}^{\>{}^{\prime}}$ to optimize the ReCo-Diff as follows:
$\mathcal{L}_{RNet}=\mathbb{E}_{\mathbf{x}_{0},\mathbf{y},t,\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\lVert\mathbf{x}_{0}-\mathcal{F}_{R}(\hat{\mathbf{x}}_{0},\hat{\mathbf{c}}_{t},\gamma_{t})\rVert^{2}_{2}].\vspace{-0.15cm}$ (17)
4 Experiments
4.1 Experimental Settings
Table 1: Quantitative comparison on the LOL [39] and LOL-v2-real [47] datasets. $\uparrow$ ($\downarrow$) denotes that, larger (smaller) values suggest better quality. The best results are highlighted in bold and the second best results are in underline (Special fonts are also used in Tabs. 2 and 3).
Methods LOL LOL-v2-real
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
LIME [11] TIP’16 16.760 0.560 0.350 117.892 15.240 0.470 0.428 118.171
RetinexNet [39] BMVC’18 16.770 0.462 0.474 113.699 18.371 0.723 0.365 133.905
KinD [52] MM’19 20.870 0.799 0.207 104.632 17.544 0.669 0.375 137.346
DRBN [46] CVPR’20 19.860 0.834 0.155 75.359 20.130 0.830 0.147 60.631
Zero-DCE [9] CVPR’20 14.861 0.562 0.335 101.237 18.059 0.580 0.313 91.939
MIRNet [50] PAMI’22 24.140 0.842 0.131 69.179 20.357 0.782 0.317 49.108
SNR [43] CVPR’22 24.608 0.840 0.151 55.121 21.479 0.848 0.157 54.532
LLFlow [36] AAAI’22 24.999 0.870 0.117 65.143 26.200 0.888 0.137 69.793
PairLIE [7] CVPR’23 19.510 0.736 0.248 100.715 20.357 0.782 0.317 96.911
SMG [44] CVPR’23 23.684 0.826 0.118 58.846 24.620 0.867 0.148 78.582
FourLLIE [24] MM’23 20.222 0.766 0.250 91.793 22.340 0.847 0.051 89.334
Retinexformer [1] ICCV’23 25.153 0.843 0.131 71.148 22.794 0.839 0.171 62.439
Diff-Retinex [48] ICCV’23 21.981 0.863 0.048 47.851 - - - -
DiffLL [14] SIGGRAPH Asia’23 26.336 0.845 0.217 48.114 28.857 0.876 0.207 45.359
PyDiff [57] IJCAI’23 27.088 0.875 0.111 56.215 27.236 0.869 0.211 57.832
ReCo-Diff (Ours) 27.626 0.884 0.090 42.990 29.306 0.906 0.105 46.677
Implementation Details. Our method is trained using an NVIDIA RTX A100 GPU and the training iteration is set to 600k. We use the Adam optimizer with the momentum as $(0.9,0.999)$. The initial learning rate is set to $1\!\times\!10^{-4}$ and decays by a factor of 0.5 after every $1\!\times\!10^{5}$ iterations. The batch size and patch size are set to 8 and 256$\times$256, respectively. We followed the similar architecture of [32] as our denoising U-Net $\bm{\epsilon}_{\theta}$. We used 2000 diffusion steps T for training. In the inference stage, we follow the DDIM [33] and set 8 steps for enhancement.
Datasets. We evaluate the proposed method on several datasets from various scenes, including LOL [39], LOL-v2 [47], and UHD-LL [24]. The LOL dataset [39] includes 485 low/normal-light image pairs for training and 15 pairs for testing. The LOL-v2 dataset [47] can be divided into two parts, the real part is an extension of LOL with 689/100 pairs and the synthetic part includes 900/100 pairs for training and testing. The UHD-LL dataset [24] includes 2000 pairs for training and 150 pairs for testing.
Metrics. To conduct comparison, we employ two full-reference metrics peak signal-to-noise ratio (PSNR) and structural similarity (SSIM) [38], and two perceptual metrics learned perceptual image patch similarity (LPIPS) [51], Fréchet Inception Distance (FID) [12] for evaluation.
Compared Methods. We compare with a rich collection of state-of-the-art methods, including LIME [11], RetinexNet [39], KinD [52], DRBN [46], Zero-DCE [9], EnlightGAN [15], MIRNet [50], SNR [43], LLFlow [36], PairLIE [7], SMG [44], FourLLIE [24], Retinexformer [1], UHDFour [24] and diffusion-based methods including Diff-Retinex [48], DiffLL [14], and PyDiff [57].
4.2 Quantitative Evaluation
Table 2: Quantitative comparison on the LOL-v2-synthetic [47].
Methods LOL-v2-synthetic
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
LIME [11] TIP’16 16.880 0.758 0.104 -
RetinexNet [39] BMVC’18 16.551 0.652 0.379 98.843
KinD [52] MM’19 18.956 0.801 0.262 89.156
DRBN [46] CVPR’20 21.687 0.825 0.174 52.972
Zero-DCE [9] CVPR’20 17.756 0.814 0.168 49.239
MIRNet [50] PAMI’22 21.941 0.876 0.112 38.775
SNR [43] CVPR’22 24.130 0.927 0.032 23.971
LLFlow [36] AAAI’22 23.692 0.911 0.103 32.576
PairLIE [7] CVPR’23 19.074 0.794 0.230 85.209
SMG [44] CVPR’23 25.618 0.905 0.053 23.210
FourLLIE [24] MM’23 24.649 0.919 0.039 26.351
RetinexFormer [1] ICCV’23 25.670 0.928 0.059 22.781
DiffLL [14] SIGGRAPH Asia’23 25.456 0.896 0.102 43.670
PyDiff [57] IJCAI’23 26.038 0.922 0.072 26.854
ReCo-Diff (Ours) 26.701 0.935 0.057 22.755
Figure 4: Visual comparison of our ReCo-Diff and the compared LLIE methods on the LOL dataset.
Figure 5: Visual comparison of our ReCo-Diff and the compared LLIE methods on the UHD-LL dataset.
Tabs. 1, 2 and 3 show the comparisons on LOL, LOL-v2, and UHD-LL. It is clear that our ReCo-Diff achieves consistent and significant performance gain over all competing methods. Specifically, our method provides significant improvement of 0.638dB/0.449dB on LOL/LOL-v2 datasets respectively, establishing the new state-of-the-art with PSNR values of 27.626dB/29.306dB. Furthermore, our method achieves similar performance in terms of SSIM, yielding the best values of 0.884/0.906. On the LOL-v2-synthetic test set, our ReCo-Diff improves the PSNR and SSIM by 0.663 and 0.013 than the second-best method. On UHD-LL, we also achieve considerable performance on PSNR and SSIM of 1.121dB and 0.012, proving our capability of generalization. Although our ReCo-Diff obtains several second-best LPIPS and FID values, it outperforms other diffusion-based LLIE methods in most cases and achieves the best FID on LOL, which indicates the overall performance on perceptual metrics is still competitive. Consequently, the considerable performance shows the capability of suppressing noise and preserving color and details by Retinex-based priors, which demonstrates the effectiveness of our proposed Retinex-based condition strategy in resolving LLIE task.
4.3 Qualitative Evaluation
Table 3: Quantitative comparison on the UHD-LL [24]. To conserve space, we select recent methods in 2023 while ensuring that the results are sufficient to demonstrate our superiority.
Methods UHD-LL
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
SMG [44] CVPR’23 25.852 0.869 0.248 41.647
FourLLIE [24] MM’23 22.462 0.814 0.296 61.380
UHDFour [24] ICLR’23 26.226 0.900 0.239 39.956
DiffLL [14] SIGGRAPH Asia’23 24.330 0.843 0.245 49.098
PyDiff [57] IJCAI’23 25.753 0.897 0.159 36.263
ReCo-Diff (Ours) 27.347 0.912 0.121 27.233
The qualitative evaluations on LOL and LOL-v2 are shown in Figs. 1, 4 and 5. As indicated by the visual comparisons, our ReCo-Diff shows superior enhancement capability and generates images with more pleasing perceptual quality. Specifically, in the first row of Fig. 4, previous methods fail to reconstruct the detailed textures of the blanket, while our ReCo-Diff provides rich details and mitigates the color gap. As for images of the middle row, although most methods enhance the white regions well, they produce incorrect color and artifacts in the regions with complex scenes. Notably, our method not only restores the print on the carton but preserves the color consistency. Furthermore, the bottom row exhibits that our ReCo-Diff is capable of recovering the vulnerable content vanishing in the results of other methods, which reasonably indicates the superior capacity of restoring naturalistic details. In Fig. 5, although PyDiff [57] shows a similar overall look to ours, the ReCo-Diff enhances the green cup and the red text with correct color and better saturation. Hence, our ReCo-Diff yields more visually pleasing results as compared to baselines, supporting our method’s excellent performance in quantitative evaluation. More results are provided in supplementary material.
4.4 Ablation Study
Figure 6: Visual comparison on the LOL/LOL-v2 datasets for investigating the contribution of key techniques of our ReCo-Diff.
Contribution of ANet, RNet, and Conditional UNet. As shown in Tab. 4, we conduct experiments of several ablation settings by removing different components from the framework individually. The “w/o UNet-Cond" denotes the removal of conditional layers in the UNet. Compared with all ablation settings, our full setting yields the best performance. In the case of the removal of FRCM leads to the decrease of PSNR value by an average of 0.364dB below the baseline, which demonstrates the effectiveness of introducing the multi-scale features from ANet into the refinement process. The comparison between “w/o UNet-Cond" and the baseline shows a similar degradation level, resulting in an average reduction of 0.349dB, which proves the necessity of integrating Retinex features into the denoising network. Notably, by comparing “w/o IRCM" and “w/ ALL", we observe a significant decline that surpasses the aforementioned two cases, which serves to highlight the crucial role of our IRCM owing to its well-designed residual refinement manner. Furthermore, the comparison between all cases under “w/ ANet" and “w/o ANet" exhibits an average drop of 0.272dB on PSNR, illustrating that the better Retinex-based condition provided by ANet eventually improves the visual quality of generated images.
Additionally, results in Fig. 6 depict that “w/o FRCM" and “UNet-Cond" lead to the reduction of regions with detailed textures, such as uneven surfaces and closely arranged chairs. Although “w/o IRCM" preserves details better, the lack of direct guidance of Retinex components induces the color shift and unnatural illumination. Fig. 7 shows that ANet reduces the noise (top rows) and enhances the illumination (bottom rows) in early steps, and improves the quality of Retinex-based conditions in the whole process.
Table 4: Ablation studies on LOL for investigating the contribution of key techniques both in the learning and refinement stages.
Learning Refinement PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
w/ ANet w/ ALL 27.626 0.884 0.090
w/o FRCM 27.344 0.875 0.112
w/o IRCM 27.262 0.872 0.119
w/o UNet-Cond 27.278 0.875 0.113
w/o ANet w/ ALL 27.301 0.877 0.103
w/o FRCM 26.956 0.869 0.131
w/o IRCM 26.783 0.864 0.140
w/o UNet-Cond 26.967 0.868 0.134
Table 5: Ablation studies of training settings of ANet for investigating the effects of the joint loss function.
Settings LOL LOL-v2-real
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
w/ Both 27.626 0.884 0.090 29.306 0.906 0.105
w/o SSIM 27.502 0.875 0.101 29.163 0.891 0.118
w/o Hist 27.508 0.875 0.102 29.156 0.892 0.119
w/o Both 27.446 0.872 0.106 29.061 0.885 0.124
Figure 7: Ablation study of ANet in terms of reflectance (row 1,3) and illumination (row 2,4) maps. Images in orange box shows the effects of ANet corresponding to Tab. 4, and green box shows the effects of SSIM and Hist loss corresponding to Tab. 5.
Joint Losses of ANet. Tab. 5 lists the results of different settings for training ANet. The “w/ Both" and “w/o Both" denote that using the original $\mathcal{L}_{ja}$ and $\mathcal{L}_{je}$ and removing both SSIM loss and Hist loss, respectively. First, “w/ Both" reasonably achieves the best performance compared with all ablation settings. The settings of “w/o SSIM" and “w/o Hist" show degradation on all metrics, demonstrating the effects of SSIM loss and Hist loss respectively. Comparing “w/ Both" and “w/o Both" shows a favorable gain of 0.210dB and illustrates the necessity of using the constraint of both the structure and histogram. Furthermore, we can see the effects of SSIM loss and Hist loss by comparing Retinex output at step 8 in Fig. 7. The reflectance and illumination maps in the green box contain more artifacts and unsatisfactory brightness distribution. More ablation studies and analysis are provided in supplementary material.
5 Conclusion
In this paper, we propose an LLIE-specific diffusion model with a Retinex-based condition strategy, ReCo-Diff, for low-light image enhancement. Our ReCo-Diff combines Retinex theory and diffusion model in two stages: Retinex-based condition learning and conditional refinement. In the learning stage, we utilize DNet to obtain the initial priors and then provide robust Retinex-based conditions by ANet. In the refinement stage, the prepared conditions are integrated into both UNet and RNet through conditional layers and RCMs to control the diffusion process and produce better color and details. Extensive experiments demonstrate that our ReCo-Diff outperforms state-of-the-art methods on representative benchmarks. We also hope to encourage more attention to develop DMs incorporating intrinsic characteristics of low-level vision.